# Kirchhoff Meets Johnson: In Pursuit of Unconditionally Secure Communication

### v2

@@ IV-C Release the Kraken: Security at the PHY Level
The high computing power and processing time requirements of modern encryption algorithms have been the major driving force behind the emergence of PHY security solutions that have flourished in the past decade. Specifically, PHY security provides node authenticity, data confidentiality, and data integrity by leveraging the dynamic characteristics of the wireless environment [22, 23, 24]. Despite the solid security solutions that PHY security offers, considering it as a quantum-resistant solution would be a naive assumption. This can be attributed to the fact that PHY security generally relies on optimistic assumptions concerning the availability of perfect channel state information (including that of the eavesdropper’s), multiple antennas, rich scattering environments, rapidly varying channels, infinite transmission power, and better channel conditions of legitimate receivers [25]. In other words, despite its solid theoretical appeal, PHY security often achieves only conditional security in many practical scenarios, meaning that Alice and Bob must have an advantage over Eve within the PHY itself. This limitation is one of the factors hindering its widespread use in current networks. Secret-key generation emerges as one of the promising applications of PHY security. In this context, Alice and Bob can leverage the wireless channel to generate correlated random sequences and subsequently use a public authenticated and error-free feedback channel to agree upon a secret key [26]. In the source-type model, terminals observe the correlated output of the randomness source to generate keys; however, challenges remain in implementing this in practice [17]. In the channel-type model, one terminal transmits random symbols to other terminals using a broadcast channel, which opens up new opportunities. Nevertheless, the security of both models can be compromised under active attacks, such as jamming and channel manipulation. Moreover, most of the available PHY key generation schemes require a noiseless public discussion channel with unlimited capacity. In light of the above discussion, secret key generation with user-generated randomness, illustrated in Fig. 5, stands out as a promising direction [26]. Here, using two transmission intervals, a shared secret key ($K_{a}=K_{b}$) is generated considering both user-generated and channel-oriented randomness and channel reciprocity. In simple terms, Alice and Bob multiply their received signals carrying the random messages of their partner by the locally generated random signals to compose a shared randomness. It is worth emphasizing the conceptual similarity between this solution and the DHM key exchange scheme.
@@
The use of AN-based systems introduces another new dimension to PHY security. In this approach, the transmitter allocates a portion of its available power to generate AN signals, strategically degrading only the eavesdropper’s channel and thereby ensuring communication security [27, 28]. Despite its significant potential, most of the AN-based systems in the literature utilize MIMO antennas to nullify the impact of AN on the legitimate receiver, essentially diminishing the signal-to-noise ratio (SNR) experienced by Eve [29]. In essence, the primary goal of AN-based systems is to reduce Eve’s SNR, not to use the AN itself for secrecy. However, assuring that Eve’s BER approaches $0.5$ when the AN power is constrained is challenging, as some confidential information might inadvertently still reach her. These challenges contribute to the conditional security nature of AN-based systems. In conclusion, while public key cryptography is widely used, it offers only conditional (computational) security and could be compromised by quantum computers within a feasible timeframe. PHY security is a complementary solution to address the computational demands and latency constraints associated with higher-level cryptographic techniques. However, it is important to note that the security of PHY solutions relies on external factors. Given these considerations, our attention naturally shifts towards physics-based key distribution systems like KLJN and QKD as we seek to attain the pinnacle of unconditional security.
@@ VI Clash Of the Titans: KLJN vs QKD
The KLJN scheme stands out as a potential alternative to QKD, using statistical-physical techniques for the secure distribution of shared randomness. Nevertheless, despite our fancy title, comparing KLJN and QKD is the comparison of apples and oranges since both schemes have entirely different architectures and characteristics. Furthermore, both designs have pros and cons and might suit different applications and users. For instance, KLJN has a much lower deployment cost and might be more suitable for low-rate ($\sim\!100$ bit/sec) key distribution applications at moderate distances ($2-20$ km) or even shorter ranges, such as inside data centers and computers [32]. Specifically, a bandwidth-distance product of $2\times 10^{6}$ Hz-meter is reported for the KLJN scheme due to the distance limitations imposed by the Kirchhoff law [6]. On the other hand, QKD systems have a significantly higher cost (\$$\sim\$100$K) due to their sophisticated hardware, however, off-the-shelf QKD products [^3] A notable commercial off-the-shelf product is the 4th generation Clavis XG QKD system from ID Quantique developed in May 2022 [33]. can support data rates of more than $100$ kbits/sec over distances higher than $100$ km. As a result, instead of putting them into competition, we have to assess their potential for specific use cases. In this sense, this article aims not to enter the ongoing debates between QKD and KLJN supporters from the past decade. Instead, it aims to shed light on their operating principles and unconditional security aspects with their potential pros and cons. Despite its undeniable potential, the literature on KLJN is relatively limited due to its multidisciplinary nature. In this context, most of the studies in the literature focus on different types of KLJN attacks and corresponding defense mechanisms. A limited number of enhanced KLJN designs and performance analyses have also been reported recently. The real-time performance of the KLJN system is investigated by practical experiments at several distances and data rates for the first time in [32] using external noise generators. In this study, a noticeable BER of $0.02\%$ is obtained through practical experiments. Wire resistance effect on the noise voltage and current, which might be inevitable in practical setups, is investigated in [34]. Here, the information leak to Eve stemming from non-zero wire resistance is reported to be not significant. It is worth noting that Gaussian-distributed noise signals (either natural background or externally generated ones) must be used in KLJN systems for the secure key exchange, as mathematically demonstrated in [35]. Furthermore, the use of artificial noise generators in practical setups allows for the generation of strong signals (noise temperatures corresponding beyond a trillion Kelvin) and the control of temperature with astronomically high accuracy [8]. An inspiring generalized KLJN scheme is introduced in [36] using arbitrary (four different) resistors. However, compared to the original KLJN scheme, this generalized scheme is shown to be less secure in the presence of passive attacks utilizing the different line resistances for $01$ and $10$ cases [37]. The theoretical BEP performance of the KLJN scheme is investigated for the first time in [38], and an exponentially decaying error probability is obtained concerning the bit duration, which might be very promising for reliable bit exchange. The joint use of voltage and current samples (measurements) is considered in the follow-up study of [39] to further reduce the BEP of KLJN systems by using more sophisticated protocols. However, both of these studies considered only the performance of securely exchanged bits, assuming that the protocol discards the insecurely detected ones and certain theoretical approximations are used to derive the theoretical BEP. In [40], using statistical hypothesis testing, a new and systematic approach is used to assess the effect of relative noise variance ratios on the BEP performance of the KLJN system. Promising analog KLJN variants such as random-resistor KLJN and random-resistor random-temperature KLJN are introduced to increase the key rate and improve the immunity against earlier attacks [41]. Recently, a comprehensive framework has been presented in [3] on the BEP optimization and the receiver system design of the KLJN scheme. In this study, not only guidelines are presented for the KLJN system designer but also two new KLJN detectors utilizing joint and alternating voltage and current samples are proposed to further improve its BEP.
@@
A variety of attack types have been proposed for the KLJN scheme in the past decade: passive time-correlation attacks [42], directional wave measurement attacks [43], transient wave attacks [44], cable resistance and capacitance based attacks [45, 46], current injection attacks [47], random number generation attacks [48], and many more. We note that some of these attacks are optimistic and might not work in practice, while others might work but have efficient defence mechanisms. While a detailed coverage of these important attacks, as well as their defense mechanisms, are beyond the scope of this review article, interested readers are referred to [8] and the references therein for a detailed investigation of these. Through the authenticated exchange of Alice and Bob’s current and voltage measurement data, a general defense method can be provided against these active (invasive) attacks, which inject or extract energy from the line [47, 7]. Comparing the currents and voltages at the two ends also eliminates the man-in-the-middle (MITM) attack for KLJN; however, this requires an authenticated and classical communication channel between Alice and Bob. In this context, QKD and KLJN schemes show similarities in their operation to maintain their integrity against specific attacks. On the contrary, academic, commercial, and governmental interest in QKD systems has flourished remarkably in the past decade. QKD systems have rapidly expanded in recent years, with numerous new commercial products and terrestrial networks being established. In the following, we present a concise overview of the QKD literature, primarily emphasizing communication engineering and algorithmic design aspects. The first QKD protocol was proposed by Bennett & Brassard in 1984, as briefly discussed earlier [31]. Specifically, different properties of photons, such as time, frequency, phase, and orbital angular momentum, can be employed to implement various QKD protocols along with continuous-variable QKD systems. Most commercial QKD products are also based on the popular BB84 protocol, a discrete-variable QKD system built on the no-cloning theorem of quantum physics and the irreversibility of measured quantum states. Due to the possibility of an MITM attack by Eve, the sifting phase must be authenticated by Alice and Bob. This phase generally consists of three messages between Alice and Bob. Two messages are sent by Bob, where he tells Alice the photons he was able to measure and the bases he used for the measurement of those. Alice completes the sifting phase by telling Bob the bases she used for those photons measured by Bob. However, this bit sifting with authentication requires a classical communication channel, as any unconditionally secure system. B92 protocol developed by Bennett in 1992 uses only two non-orthogonal states in Alice to represent bit 00 and bit $1$ [49]. An important contrast to the BB84 protocol is that Bob does not have to announce the choice of his bases in the post-transmission communication. In other words, bit sifting is not required in BB92 due to its clever polarization selection algorithm. The SARG04 protocol uses the same four states and the same measurements on Bob’s side as BB84, but the bit is coded in the basis rather than in the state [50]. The advantage of this protocol is that Alice does not need to reveal her choice of basis. After all these three protocols, Alice and Bob perform classical post-processing, including information reconciliation and privacy amplification, to remove the errors stemming from the channel and to reduce the correlation that remains with Eve, respectively. An overview of the rich set of QKD protocols can be found in [51].

### v3

@@ IV-C Release the Kraken: Security at the PHY Level
The high computing power and processing time requirements of modern encryption algorithms have been the major driving force behind the emergence of PHY security solutions that have flourished in the past decade. Specifically, PHY security provides node authenticity, data confidentiality, and data integrity by leveraging the dynamic characteristics of the wireless environment [22, 23, 24]. Despite the solid security solutions that PHY security offers, considering it as a quantum-resistant solution would be a naive assumption. This can be attributed to the fact that PHY security generally relies on optimistic assumptions concerning the availability of perfect channel state information (including that of the eavesdropper’s), multiple antennas, rich scattering environments, rapidly varying channels, infinite transmission power, and better channel conditions of legitimate receivers [25]. In other words, despite its solid theoretical appeal, PHY security often achieves only conditional security in many practical scenarios, meaning that Alice and Bob must have an advantage over Eve within the PHY itself. This limitation is one of the factors hindering its widespread use in current networks.
Secret-key generation emerges as one of the promising applications of PHY security. In this context, Alice and Bob can leverage the wireless channel to generate correlated random sequences and subsequently use a public authenticated and error-free feedback channel to agree upon a secret key [26]. In the source-type model, terminals observe the correlated output of the randomness source to generate keys; however, challenges remain in implementing this in practice [17]. In the channel-type model, one terminal transmits random symbols to other terminals using a broadcast channel, which opens up new opportunities. Nevertheless, the security of both models can be compromised under active attacks, such as jamming and channel manipulation. Moreover, most of the available PHY key generation schemes require a noiseless public discussion channel with unlimited capacity.
In light of the above discussion, secret key generation with user-generated randomness, illustrated in Fig. 5, stands out as a promising direction [26]. Here, using two transmission intervals, a shared secret key ($K_{a}=K_{b}$) is generated considering both user-generated and channel-oriented randomness and channel reciprocity. In simple terms, Alice and Bob multiply their received signals carrying the random messages of their partner by the locally generated random signals to compose a shared randomness. It is worth emphasizing the conceptual similarity between this solution and the DHM key exchange scheme.
@@
The use of AN-based systems introduces another new dimension to PHY security. In this approach, the transmitter allocates a portion of its available power to generate AN signals, strategically degrading only the eavesdropper’s channel and thereby ensuring communication security [27, 28]. Despite its significant potential, most of the AN-based systems in the literature utilize MIMO antennas to nullify the impact of AN on the legitimate receiver, essentially diminishing the signal-to-noise ratio (SNR) experienced by Eve [29]. In essence, the primary goal of AN-based systems is to reduce Eve’s SNR, not to use the AN itself for secrecy. However, assuring that Eve’s BER approaches $0.5$ when the AN power is constrained is challenging, as some confidential information might inadvertently still reach her. These challenges contribute to the conditional security nature of AN-based systems.
In conclusion, while public key cryptography is widely used, it offers only conditional (computational) security and could be compromised by quantum computers within a feasible timeframe. PHY security is a complementary solution to address the computational demands and latency constraints associated with higher-level cryptographic techniques. However, it is important to note that the security of PHY solutions relies on external factors. Given these considerations, our attention naturally shifts towards physics-based key distribution systems like KLJN and QKD as we seek to attain the pinnacle of unconditional security.
@@ V Quantum Key Distribution: Game Over for Codebreakers?
QKD addresses one of society’s most pressing concerns: unconditionally secure communication. Specifically, QKD systems leverage the principles of quantum mechanics to create a secret key between two distant users. More importantly, the security of this key is guaranteed by the laws of physics [17, 30]. A unique property of QKD is its users’ ability to detect the presence of a third party in between, either passive or active since measuring a quantum system disturbs it. In other words, an eavesdropper trying to gain access to the key must perform measurements, which cause detectable anomalies at QKD transmitter and receiver nodes. Lastly, it is noteworthy that while QKD cannot be used for secure data transmission, similar to KLJN, it can generate secure keys for subsequent bulk data encryption, i.e., it is a key establishment protocol.
@@
Since Bob does not know the bases Alice encoded in the photons, he can only select a basis at random to measure the incoming photon, either rectilinear or diagonal bases (rectilinear or diagonal detectors). After this step, Alice and Bob hold a bit sequence called the raw key pair. At this point, they rely on their public discussions to filter out this raw key pair. Through a public channel, Bob announces the bases he has randomly selected to measure Alice’s photons. After comparing Bob’s bases to the ones she used in photon generation, Alice responds to Bob by stating the cases in their selections coincide. Finally, Alice and Bob discard the corresponding bits for which the encoding and measurement bases are not the same, and they obtain the shared secret (sifted) key.
Let us give an example considering the QKD scheme given in Fig. 6. For the first transmission, Alice and Bob’s bases coincide (both rectilinear); as a result, Bob’s detection of bit $1$ is correct. On the contrary, for the second transmission, Bob’s randomly selected basis (diagonal) is different from that of Alice’s (rectilinear), and the bit interpretation of Bob is inconclusive. Specifically, Bob can obtain either bit 00 or bit $1$ with $50\%$ probability after detection with the wrong basis, and he discards this bit in the sifted key after public discussion of selected bases with Alice. For the transmission given in Fig. 6 with fourteen bits, the bases of Alice and Bob coincide for photons $1,4,5,6,8,9,10,12,$ and $14$, and the shared key is obtained as $110010010$ after public discussion of selected bases.
After 2000 years, is it game over for codebreakers? At different times in history, cryptographers have believed that the monoalphabetic/polyalphabetic ciphers and mechanical ciphers such as Enigma were all unbreakable; however, their claims were eventually proved wrong. This can be explained by the fact that all these ciphers were developed, assuming that they were complex enough to challenge the ingenuity and the technology of codebreakers at that time. On the other hand, quantum cryptography exploiting QKD is totally unbreakable under ideal conditions, and this claim is fundamentally different from the earlier ones. This is one of the reasons why codebreakers are transforming into physicists and engineers to understand and exploit the practical imperfections of QKD (and KLJN) systems to gain knowledge on secure keys. Here, we can ask the following questions: What are the significant disadvantages of QKD in practice? Why is post-quantum cryptography favored more than QKD for practical networks, for instance, by authorities [12]? These exciting questions remain to be answered.
@@ VI Clash Of the Titans: KLJN vs QKD
The KLJN scheme stands out as a potential alternative to QKD, using statistical-physical techniques for the secure distribution of shared randomness. Nevertheless, despite our fancy title, comparing KLJN and QKD is the comparison of apples and oranges since both schemes have entirely different architectures and characteristics. Furthermore, both designs have pros and cons and might suit different applications and users. For instance, KLJN has a much lower deployment cost and might be more suitable for low-rate ($\sim\!100$ bit/sec) key distribution applications at moderate distances ($2-20$ km) or even shorter ranges, such as inside data centers and computers [32]. Specifically, a bandwidth-distance product of $2\times 10^{6}$ Hz-meter is reported for the KLJN scheme due to the distance limitations imposed by the Kirchhoff law [6]. On the other hand, QKD systems have a significantly higher cost (\$$\sim\$100$K) due to their sophisticated hardware, however, off-the-shelf QKD products [^3] A notable commercial off-the-shelf product is the 4th generation Clavis XG QKD system from ID Quantique developed in May 2022 [33]. can support data rates of more than $100$ kbits/sec over distances higher than $100$ km. As a result, instead of putting them into competition, we have to assess their potential for specific use cases. In this sense, this article aims not to enter the ongoing debates between QKD and KLJN supporters from the past decade. Instead, it aims to shed light on their operating principles and unconditional security aspects with their potential pros and cons.
VI-A A Brief Review of KLJN Literature
Despite its undeniable potential, the literature on KLJN is relatively limited due to its multidisciplinary nature. In this context, most of the studies in the literature focus on different types of KLJN attacks and corresponding defense mechanisms. A limited number of enhanced KLJN designs and performance analyses have also been reported recently.
The real-time performance of the KLJN system is investigated by practical experiments at several distances and data rates for the first time in [32] using external noise generators. In this study, a noticeable BER of $0.02\%$ is obtained through practical experiments. Wire resistance effect on the noise voltage and current, which might be inevitable in practical setups, is investigated in [34]. Here, the information leak to Eve stemming from non-zero wire resistance is reported to be not significant. It is worth noting that Gaussian-distributed noise signals (either natural background or externally generated ones) must be used in KLJN systems for the secure key exchange, as mathematically demonstrated in [35]. Furthermore, the use of artificial noise generators in practical setups allows for the generation of strong signals (noise temperatures corresponding beyond a trillion Kelvin) and the control of temperature with astronomically high accuracy [8].
An inspiring generalized KLJN scheme is introduced in [36] using arbitrary (four different) resistors. However, compared to the original KLJN scheme, this generalized scheme is shown to be less secure in the presence of passive attacks utilizing the different line resistances for $01$ and $10$ cases [37]. The theoretical BEP performance of the KLJN scheme is investigated for the first time in [38], and an exponentially decaying error probability is obtained concerning the bit duration, which might be very promising for reliable bit exchange. The joint use of voltage and current samples (measurements) is considered in the follow-up study of [39] to further reduce the BEP of KLJN systems by using more sophisticated protocols. However, both of these studies considered only the performance of securely exchanged bits, assuming that the protocol discards the insecurely detected ones and certain theoretical approximations are used to derive the theoretical BEP. In [40], using statistical hypothesis testing, a new and systematic approach is used to assess the effect of relative noise variance ratios on the BEP performance of the KLJN system. Promising analog KLJN variants such as random-resistor KLJN and random-resistor random-temperature KLJN are introduced to increase the key rate and improve the immunity against earlier attacks [41]. Recently, a comprehensive framework has been presented in [3] on the BEP optimization and the receiver system design of the KLJN scheme. In this study, not only guidelines are presented for the KLJN system designer but also two new KLJN detectors utilizing joint and alternating voltage and current samples are proposed to further improve its BEP.
In Fig. 7, we present the timeline of some major KLJN developments and the corresponding studies from the literature reviewed in this subsection.
Figure 7: Timeline of some major KLJN developments.
A variety of attack types have been proposed for the KLJN scheme in the past decade: passive time-correlation attacks [42], directional wave measurement attacks [43], transient wave attacks [44], cable resistance and capacitance based attacks [45, 46], current injection attacks [47], random number generation attacks [48], and many more. We note that some of these attacks are optimistic and might not work in practice, while others might work but have efficient defence mechanisms. While a detailed coverage of these important attacks, as well as their defense mechanisms, are beyond the scope of this review article, interested readers are referred to [8] and the references therein for a detailed investigation of these.
Through the authenticated exchange of Alice and Bob’s current and voltage measurement data, a general defense method can be provided against these active (invasive) attacks, which inject or extract energy from the line [47, 7]. Comparing the currents and voltages at the two ends also eliminates the man-in-the-middle (MITM) attack for KLJN; however, this requires an authenticated and classical communication channel between Alice and Bob. In this context, QKD and KLJN schemes show similarities in their operation to maintain their integrity against specific attacks.
@@ VI-A A Brief Review of KLJN Literature
VI-B Background of QKD Protocols
On the contrary, academic, commercial, and governmental interest in QKD systems has flourished remarkably in the past decade. QKD systems have rapidly expanded in recent years, with numerous new commercial products and terrestrial networks being established. In the following, we present a concise overview of the QKD literature, primarily emphasizing communication engineering and algorithmic design aspects.
The first QKD protocol was proposed by Bennett & Brassard in 1984, as briefly discussed earlier [31]. Specifically, different properties of photons, such as time, frequency, phase, and orbital angular momentum, can be employed to implement various QKD protocols along with continuous-variable QKD systems. Most commercial QKD products are also based on the popular BB84 protocol, a discrete-variable QKD system built on the no-cloning theorem of quantum physics and the irreversibility of measured quantum states. Due to the possibility of an MITM attack by Eve, the sifting phase must be authenticated by Alice and Bob. This phase generally consists of three messages between Alice and Bob. Two messages are sent by Bob, where he tells Alice the photons he was able to measure and the bases he used for the measurement of those. Alice completes the sifting phase by telling Bob the bases she used for those photons measured by Bob. However, this bit sifting with authentication requires a classical communication channel, as any unconditionally secure system.
B92 protocol developed by Bennett in 1992 uses only two non-orthogonal states in Alice to represent bit 00 and bit $1$ [49]. An important contrast to the BB84 protocol is that Bob does not have to announce the choice of his bases in the post-transmission communication. In other words, bit sifting is not required in BB92 due to its clever polarization selection algorithm. The SARG04 protocol uses the same four states and the same measurements on Bob’s side as BB84, but the bit is coded in the basis rather than in the state [50]. The advantage of this protocol is that Alice does not need to reveal her choice of basis.
After all these three protocols, Alice and Bob perform classical post-processing, including information reconciliation and privacy amplification, to remove the errors stemming from the channel and to reduce the correlation that remains with Eve, respectively. An overview of the rich set of QKD protocols can be found in [51].
@@ VI-B Background of QKD Protocols
Two prevalent QKD attacks are intercept-resend (IR) and photon number splitting (PNS) attacks [52, 17]. In the former, Eve intercepts the quantum signal sent by Alice and measures it. Based on the measurement result, she prepares a new quantum signal in the same measured state to be sent to Bob. Due to quantum weirdness, Eve can only detect correct states with a probability of $1/2$. In return, the overall error probability between Alice and Bob becomes $1/4$, which is relatively high and can be easily noticed by them. However, in the more sophisticated PNS attack, Eve exploits the fact that Alice transmits multiple photons for each bit. Specifically, Alice’s coherent source emits photons according to the Poisson distribution with a typical average photon number of $0.1-0.8$ (Poisson parameter $\alpha$) for weak laser pulses [30]. Despite its rare occurrence probability, in this case, Eve employs a beamsplitter to take one of the photons and pass the rest to Bob. To circumvent this, Alice can employ a decoy-state-based protocol. In this protocol, to represent the signal and decoy states, Alice transmits quantum states with different mean photon numbers to confuse Eve [53]. While the literature contains a rich set of other attacks involving heavy quantum physics and numerous debates about the practical vulnerabilities of QKD systems, we conclude our discussion at this point.
VI-C KLJN vs QKD
Table I presents a high-level comparison of KLJN and QKD systems considering their major key performance indicators (KPIs), such as key rate, range, complexity, cost, and BER. Here, we use high, medium, and low labels based on specific KPIs for each system, and red and green colors are respectively used for negative and positive aspects of each system for the corresponding KPI. For instance, QKD outperforms KLJN regarding key rate, benefiting from not requiring a circuit loop between two users. Nevertheless, sophisticated attacks utilizing photon groups can limit the key rate of practical QKD systems. However, QKD and KLJN systems have different power scaling laws and the target range is critical in their comparison. We note that the complexity of KLJN can be as high as QKD when fully defended but its cost can be very low if integrated on a chip. In the same table, we also list the corresponding major protocols associated with these two physics-based key distribution systems. We conclude that both designs can offer unconditional security by exploiting the laws of physics under certain conditions. However, the choice between these two key distribution approaches should be made carefully, depending on the requirements of the target applications. Evaluating the performance of two concepts through computer simulations and practical experiments is also a promising future research direction. At this point, we also pose the following interesting question: can these two different key distribution concepts merged to create an ultimate key distribution system?
Finally, it is worth noting that both KLJN and QKD systems are affected by real-world conditions and can be used cautiously by system designers in the presence of certain effects. For instance, the thermal equilibrium of the two resistors of the KLJN system should be preserved to guarantee perfect security and eliminate cable resistance-based attacks [8]. Furthermore, in practical setups, wire resistance should be much less than Alice and Bob’s resistances, and the wire capacitance should not affect the loop impedance. Proper filtering and advanced detectors might also be necessary to reach the theoretical key rates of KLJN systems and to overcome the deviations caused by environmental factors. Furthermore, there might be an information leak in non-ideal practical setups, while unconditional security can still be preserved by various techniques such as privacy amplification. On the other hand, practical QKD systems also suffer from short range, low bit rate, and high cost. Furthermore, generating strictly ideal single photon packages and avoiding detector noise is impossible for QKD systems under practical constraints.
To conclude, both KLJN and QKD systems rely on specific laws of physics to create secure keys; however, they might require additional defense mechanisms in practice to preserve unconditional security. If implemented at the chip level, the low cost of KLJN systems makes them a potential candidate for short-range critical applications in which system designers can create effective Kirchhoff loops. On the other hand, QKD systems require dedicated hardware, have a much higher cost, and can be suitable for more advanced and long-range key distribution applications.